OpenStudy (ray10):

Find all fourth roots of \[-8-8\sqrt{3}i\] ? I come to r=16, \[\theta \] = \[\frac{ \Pi }{ 3}\] , then get lost

terenzreignz (terenzreignz):

That is one big pi. Use \pi instead of \Pi lol

OpenStudy (ray10):

haha ah thank you! I shall :)

OpenStudy (ray10):

\[\frac{ \pi }{ 3 }\]

terenzreignz (terenzreignz):

Isn't that better? :)

OpenStudy (ray10):

yes it certainly is thanks :) didn't know about that. Do you think you can help me with this question? :)

terenzreignz (terenzreignz):

Can I indeed...

terenzreignz (terenzreignz):

True enough, the modulus r is 16, but tell me how you acquired \(\Large \frac \pi 3\) for your argument.

terenzreignz (terenzreignz):

My guess is that you naïvely took \[\Large \tan^{-1}\left(\frac{-8\sqrt3}{-8}\right)\]

OpenStudy (ray10):

\[\tan^{-1} (\frac{ -8\sqrt{3} }{ -8 }) = \frac{ \pi }{ 3 }\]

OpenStudy (ray10):

the other way I looked at it was, without the negative in front of the \[-8\sqrt{3}\], I thought that was wrong

terenzreignz (terenzreignz):

Unacceptable. :) You do realise that \(\large \tan^{-1}\) is not the most reliable of functions. Your logic is far more dependable in these matters.

terenzreignz (terenzreignz):

Now, you know that the tangent of the argument is \(\large \sqrt3\)

terenzreignz (terenzreignz):

A positive, yet your real and imaginary parts are negative, implying naturally that the angle must be in the third quadrant of the Argand Diagram or complex plane.

OpenStudy (ray10):

hmm alright, I can understand where you are coming from :) yes I see the logic behind it being \[\sqrt{3}\]

terenzreignz (terenzreignz):

The tangent of the angle IS \(\Large \sqrt 3\) and this comes from the elementary division \[\Large \frac{-8\sqrt3}{-8}=\frac{\cancel{-8}\sqrt3}{\cancel{-8}}\]

OpenStudy (ray10):

therefore \[\theta \] = \[\sqrt{3}\] :)

terenzreignz (terenzreignz):

Most definitely not. \[\Large \color{red}{\tan}\theta = \sqrt3\] Do not confuse these things.

OpenStudy (ray10):

oh right because we are still at the point \[\tan \theta = \frac{ -8\sqrt{3} }{ -8 } = \sqrt{3}\]

terenzreignz (terenzreignz):

yes. now do you see a problem?

OpenStudy (ray10):

hmmm sort of, now we must solve for \[\theta\] ?

OpenStudy (ray10):

but in the third quadrant ?

terenzreignz (terenzreignz):

It would appear this is not the first time you've asked this question.

OpenStudy (ray10):

yes it's not the first time, I seem to get lost at this point often :/

terenzreignz (terenzreignz):

Strange... So what should be \(\large \theta\)?

OpenStudy (ray10):

\[\theta\] would equal \[\frac{ \pi }{ 3}\]

OpenStudy (ray10):

but would be required in the third quadrant right?

terenzreignz (terenzreignz):

I'm almost compelled to mimic Pan and say "I already implied that \(\Large \frac \pi 3\) is wrong, but something tells me you're about to arrive at the correct angle anyway...

OpenStudy (ray10):

so therefore be \[\frac{ 4 \pi }{ 3 }\] :D

terenzreignz (terenzreignz):

You're not just reading that previous thread, are you? You DO understand why that's the correct angle, right?

OpenStudy (ray10):

because we are looking at the third quadrant, therefore we are required to find the equal such angle in the third quadrant, being that of \[\pi + \frac{ \pi }{ 3 }\]

OpenStudy (ray10):

no I'm not reading that

terenzreignz (terenzreignz):

Not the explanation I was looking for, but in the assumption that you know what you're doing, then okay, that's good enough.

OpenStudy (ray10):

what would be the explanation you would say? :)

terenzreignz (terenzreignz):

Well, now, there you have it, you have the modulus r=16 and the argument \(\large \theta = \frac{4\pi}3\)

terenzreignz (terenzreignz):

The explanation I'd say is that we're looking for the angle whose tangent is \(\large \sqrt3\) and lies in the third quadrant.

OpenStudy (ray10):

Yes I see how that explains it neatly. so now we have \[r=16\] \[\theta = \frac{ 4\pi }{ 3 }\] and \[16[cis(\frac{ 4\pi }{ 3 })]\]

terenzreignz (terenzreignz):

yes indeed. Now find the fourth roots

OpenStudy (ray10):

that is where I have trouble with, Pan mentioned to add \[\frac{ \pi }{ 2 }\] but I thought we are meant to add \[2\pi \]

terenzreignz (terenzreignz):

You add \(2\pi\) to the original angle but and THEN you multiply the new angle by \(\Large \frac14\) to get fourth roots... in effect, you're just adding \(\Large \frac{2\pi}{4}\) to the FOURTH ROOT angle, or \(\Large \frac\pi2\)

terenzreignz (terenzreignz):

That might have been a little tough to digest. You either add \(2\pi\) here: \[\Large \left\{16\left[\text{cis}\left(\frac{ 4\pi }{ 3 }+\color{red}{2\pi}\right)\right]\right\}^\frac{1}{4}\] OR you add \(\Large \frac\pi 2\) here: \[\Large \left.2\left[\text{cis}\left(\frac{ \pi }{ 3 }+\color{red}{\frac\pi2}\right)\right]\right.\]

terenzreignz (terenzreignz):

They mean the same thing.

OpenStudy (ray10):

1. \[\frac{ 4\pi }{ 3 } +2\pi\] = \[\frac{ 10\pi }{ 3 }\] 2. \[\frac{ 10\pi }{ 3 } \times \frac{ 1 }{ 4 } = \frac{ 5\pi }{ 6 }\] 3. \[\frac{ 5\pi }{ 6 } \times ?\]

OpenStudy (ray10):

Pan showed me how to do this, but waking up today, I don't get the part where we find each root after the first one :S

terenzreignz (terenzreignz):

wait...

terenzreignz (terenzreignz):

\[\Large \frac{5\pi}{6}\] is right.

terenzreignz (terenzreignz):

Okay, assuming you found the first root, which happens to be \[\Large 2\text{cis}\left(\frac{\pi}{3}\right)\]

OpenStudy (ray10):

I get lost after that :P I multiply by \[\frac{ 1 }{ 4 }\]

terenzreignz (terenzreignz):

Okay, relax.... just take a deep breath and ready yourself...

OpenStudy (ray10):

I get lost after that :P I multiply by \[\frac{ 1 }{ 4 }\]

terenzreignz (terenzreignz):

We have one fourth root, right? \[\Large 2\text{cis}\left(\frac{\pi}{3}\right)\]

terenzreignz (terenzreignz):

I'm going to ask you to do something a little crazy, but it might just ease up your load for now.

OpenStudy (ray10):

yes that is correct :) okay hit me! :D

terenzreignz (terenzreignz):

Forget adding 2pi. As soon as you found one fourth root, such as this one \[\Large 2\text{cis}\left(\frac{\pi}{3}\right)\] then just keep adding 2pi/4 (4 in this case because we're looking for fourth roots)

terenzreignz (terenzreignz):

just keep adding 2pi/4 to the angle until you have a total of FOUR fourth roots. (if we were looking for fifth roots, then you'd have to get one fifth root, and then keep adding 2pi/5 until you get a total of FIVE fifth roots)

terenzreignz (terenzreignz):

So, root number 1: \[\Large 2\text{cis}\left(\frac{\pi}{3}\right)\] Root number 2: \[\Large 2\text{cis}\left(\frac{\pi}{3}+\frac{\pi}{2}\right)\]

OpenStudy (ray10):

O.o, it worked! And I really understood that O.o

terenzreignz (terenzreignz):

I make a point of making sure I'm understood, at least, to the best of my ability. Don't tell my you understood it... SHOW me.

terenzreignz (terenzreignz):

me*

OpenStudy (ray10):

\[\frac{ \pi }{ 3 }, (\frac{ \pi }{ 3 }+\frac{ \pi }{ 2 })=\frac{ 5\pi }{ 6 },(\frac{ 5\pi }{ 6 }+\frac{ \pi }{ 2 })= \frac{ 4\pi }{ 3 }, (\frac{ 4\pi }{ 3 }+\frac{ \pi }{ 2 })=\frac{ 11\pi }{ 6 } \]

terenzreignz (terenzreignz):

Again, I'm trusting that you didn't get this from simply reading Pan's work :)

OpenStudy (ray10):

I hope I am right this time :S

terenzreignz (terenzreignz):

Note that those, in and of themselves aren't the answers yet.

OpenStudy (ray10):

noo I didn't read his work, you said that I must add \[\frac{ 2\pi }{ 4 }\] to each new angle

terenzreignz (terenzreignz):

But \[\Large 2\text{cis}\left(\frac{\pi}{3}\right)\]\[\Large 2\text{cis}\left(\frac{5\pi}{6}\right)\]\[\Large 2\text{cis}\left(\frac{4\pi}{3}\right)\]\[\Large 2\text{cis}\left(\frac{11\pi}{6}\right)\]

terenzreignz (terenzreignz):

I'll leave you to work these out, I have confidence in you :)

OpenStudy (ray10):

you said a little while ago, "I'm going to get you to do something crazy" I understand the ease of use, but what other way is standard to be used? :)

OpenStudy (ray10):

Thank you @terenzreignz !!!

terenzreignz (terenzreignz):

Actually, this way is used so often, this may well BE the standard way.

terenzreignz (terenzreignz):

I was just worried that you'll love the shortcut so much (the secret of the roots?) that you'll forget why it works.

OpenStudy (ray10):

oh woah, that is a relief to hear! :) oh no not at all, I asked again how to do this question because I want to have the confidence to teach this myself :)

terenzreignz (terenzreignz):

You're teaching this?

OpenStudy (ray10):

not teaching as teacher, but say if another colleague of mine is stuck, I want to be able to help them while knowing the foundation behind it :)

terenzreignz (terenzreignz):

Well because while adding 2\pi to the original angle: \[\Large 16\text{cis}\left(\frac{4\pi}{3}+2\pi\right)\] doesn't change it, it DOES give you a new fourth root (or any nth root for that matter) \[\Large 2\text{cis}\left[\frac14\left(\frac{4\pi}{3}+2\pi\right)\right]= 2\text{cis}\left[\frac{\pi}{3}+\frac\pi2\right]\]

terenzreignz (terenzreignz):

Understood?

OpenStudy (ray10):

ahh I see, so the way you asked me to do it just then is another way, and I prefer the way you taught me. This way I must multiply by \[\frac{ 1 }{ 4 }\] after adding \[2\pi \] ?

terenzreignz (terenzreignz):

Yes, that's true. That's exactly the same way. How observant of you :)

terenzreignz (terenzreignz):

Don't you just feel smarter? :)

OpenStudy (ray10):

I sure do!! It's always great knowing why you're doing something than just knowing how :D

terenzreignz (terenzreignz):

I'm sure. But you were still outplayed by a 14-year-old :P (whom I taught) By the way, it's TJ (for Terence Jason) to you ^_^

OpenStudy (ray10):

I know xD Pan is incredibly good! :D I'm wondering how old you are now :O TJ, very nice to meet you, and you certainly helped me out really well!

terenzreignz (terenzreignz):

I'm 17, as my profile states ^_^ I turned 17 two months ago. Well, I have to go to school, so... I'll be seeing you, I guess ^_^ ---------------------------------------- Terence out

OpenStudy (ray10):

talk soon! Have a good time in school ^_^ thanks heaps again!